Really need help on these.
Find all the solution's for each triangle.Round to the nearest 10th.
1.a=125,A=25,b=150
2.a=32,c=20,A=112

Answer by stanbon(75887)

Find all the solution's for each triangle.Round to the nearest 10th.
1.a=125,A=25,b=150
-------------------
I'll do the 1st one.
Since you have an angle and the side opposite, use the Law of Sines.
--------------
Sin(B)/b = sin(A)/a
sin(B)/150 = sin(25)/125
sin(B) = 150[sin(25)/125]
sin(B) = 0.50714...
B = arcsin(0.50714) = 30.5 degrees
--------------------------------------
Then C = 180 - (30.5+25)= 124.5 degrees
-------------------------
Then c/sin(C) = a/sin(A)
c = sin(124.5)[125/sin(25)]
c = 243.7
=================
Cheers,
Stan H.